# Scaling of highly excited Schrödinger-Poisson eigenstates and universality of their rotation curves

### v1

@@ Abstract
This work provides a comprehensive numerical characterization of the excited spherically symmetric stationary states of the Schrödinger-Poisson problem. The analysis is conducted with reference to dark matter distributions around galaxies; the squared modulus of the eigenfunction represents the dark matter density distribution, and an associated eigenvelocity models the experimental galactic rotation curves. This paper aims to establish a connection between the fundamental structure of these eigenfunctions and eigenvelocities and the physical properties of the system.
Through numerical simulations of highly excited eigenstates, novel heuristic laws are proposed, which describe how their fundamental features scale with the excitation index $n$. Key characteristics of the eigenfunctions include: the effective support, which exhibits a parabolic dependence on the excitation index; the distances between adjacent nodes, whose pattern varies regularly with $n$; and the oscillation amplitude, which follows a power law with an exponent approaching $-1$ for large $n$, consistent with astrophysical observations. The eigenvelocities exhibit a mid-range oscillatory region with an average linear trend, whose slope approaches zero in the large $n$ limit, consistent with the physical expectation of a flat plateau for the rotation curves. Additionally, we derive heuristic scaling relationships with the excitation index $n$, revealing a universal behavior in the predicted rotation curves.
@@
Schrödinger-Poisson , Scaling , Excited stationary states , Rotation curves , Universality , Plateau
@@
[1] organization=Department of Mathematics “Tullio Levi-Civita”, University of Padova, addressline=Via Trieste, 63, city=Padova, postcode=35121, country=Italy
@@
[3] organization=Padua Quantum Technologies Research Center (QTech), addressline=Via Gradenigo, 6/A, city=Padova, postcode=35121, country=Italy
@@ 1 Introduction
The Schrödinger-Poisson model is currently the leading description of dark matter distributions around galaxies [1, 2]. Based on the assumption of ultralight particles [2, 3, 4, 5, 6], it describes the evolution of a matter field $\psi(t,x)$ via the Schrödinger equation, where $|\psi|^{2}(t,x)$ represents the dark matter particle density distribution. The particles are subjected to self-gravity, which is modeled through a Poisson potential $\phi(t,x)$, with the particle density $|\psi|^{2}(t,x)$ acting as the source. In dimensionless form, the system reads:
@@
Current dark matter distributions are generally considered to have reached a stationary state. Consequently, they are typically modeled using the Schrödinger-Poisson stationary states (see [3, 5]) – solutions to system (1) of the form $(e^{i\varepsilon t}f(x),\phi(x))$. The state-of-the-art primarily focuses on radially symmetric stationary states, which comprise an infinite, discrete family $\{\varepsilon_{n},\,f_{n}(r),\,\phi_{n}(r)\}_{n=0}^{\infty}$, with positive eigenvalues $\varepsilon_{n}>0$, real eigenfunctions $f_{n}(r)$ and excitation index $n\in\mathbb{N}$ (see [7]). As suggested by Sin [5], and as shown in Figure 1, the $n$-th excited stationary state represents dark matter distributions as $n+1$ concentric shells, separated by voids and exhibiting decreasing density outwards. A rigorous characterization of this structure and its dependence on the excitation index $n$ remains incomplete, which this paper aims to address.
Figure 1: Three-dimensional section of the dark matter distribution, as predicted by the $8$-th stationary state $f_{8}(r)$. For visualization purposes, the modulus $|f_{8}|(r)$ is shown rather than the actual predicted matter density $|f_{8}|^{2}(r)$. The structure exhibits $9$ shells of decreasing density separated by $8$ voids (shown in dark blue).
Given a predicted dark matter distribution $f_{n}^{2}(r)$ , one can calculate the tangential velocity of a test object in circular orbit at radius $r$ from the distribution’s center. This calculation assumes the velocity is determined by the gravitational force exerted by matter within a sphere of radius $r$ (see Sin [5]):
@@
This prediction plays a crucial role, as it can be compared with experimental rotation curves of spiral galaxies [8, 9, 10, 11, 12], providing both model validation and deeper physical insights into the problem. Thus, we aim to provide a comprehensive, quantitative description of the eigenvelocities (2) arising from the Schrödinger-Poisson model, characterizing their key features and relating them to established physical expectations.
@@ 1.1 Comparison with previous works and main results
However, a comprehensive analysis of how properties scale with the excitation index $n$ remains lacking. To address this gap, we propose quantitative, heuristic laws derived from numerical simulations of highly excited stationary states ($n\leq 80$), aiming to provide foundations for future analytical and physical studies. Specifically, we analyze the eigenfunctions’ structure, characterizing them as oscillatory functions modulated by a power law. The radial position of the last oscillation, which approximately defines the eigenfunction support, exhibits a clear parabolic dependence on $n$, characteristic of Keplerian problems like the Bohr model for the hydrogen atom (see e.g. [22]). The spacing between adjacent nodes increases with the radial position, following a consistent pattern across increasing $n$ values – showing modest increases at small radii and more pronounced increases at larger radii. This behavior constitutes a key correction to the $\sin{(r)}r^{-1}$ approximation, which fails to account for irregular nodal spacing. Finally, the power law governing amplitude modulation is characterized by fitting local extrema at small radial positions, with deviations observed near the last extremum. The power law exponent approaches $-1$ asymptotically as $n$ increases, recovering the $r^{-1}$ approximation in the large $n$ limit.
From a physical perspective, rotation curves serve as the primary experimental reference for comparing with analytical predictions. Observed rotation curves typically show mid-range oscillatory behavior with an approximately linear trend, forming a plateau. This plateau generally exhibits a near-flat slope, though both positive and negative slopes have also been documented (see e.g. [23, 24, 8]). Moreover, physical intuition suggests that galactic-scale dark matter clusters, which generate these rotation curves, might exhibit some kind of universal behavior, as first noted by Rubin et al. [25], Burstein and Rubin [26].
The numerical results presented in this paper align well with these physical observations. In particular, the large-$n$ approximation $f_{n}(r)\sim r^{-1}$, derived from the local extrema fit, suggests that the corresponding eigenvelocity (2) remains approximately constant, consistent with an expected flat plateau. Finally, we find that the velocity at the last local oscillation exhibits a clear dependence on its radial position, which itself varies with $n$. This relationship reveals appropriate scaling laws under which all predicted eigenvelocities converge to a similar shape, independent of $n$, exhibiting some kind of universal behavior in the predicted rotation curves and confirming that our model effectively captures expected physical properties.
@@ 2 Model
The Schrödinger-Poisson model (1) describes the dynamics of a particle density $|\psi|^{2}(t,x)$ under self-gravity. The model couples a Schrödinger equation for the matter field $\psi(t,x)$ with a Poisson equation for the potential $\phi(t,x)$. Its stationary states are expressed in the form $(e^{i\varepsilon t}f(x),\phi(x))$, where the phase oscillation in time naturally emerges from the norm conservation law $\partial_{t}(\int_{\mathbb{R}^{3}}|\psi(t,x)|^{2}\mathop{}\!\mathrm{d}^{3}x)=0$, with $\varepsilon$ representing the associated energy parameter. The spatial component $f(x)$ of the matter field can be assumed real without loss of generality.
@@
where equation (4a) is known as the Choquard equation and represents a nonlinear eigenvalue problem, with eigenvalue $\varepsilon$ and eigenfunction $f(x)$. Following the prevailing approach in the literature, we focus on spherically symmetric solutions [^1] The existence of non-spherical solutions of the stationary Schrödinger-Poisson system (4c) is still an unsolved problem (see Hui et al. [2], Appendix B)., which form an infinite, discrete family $\{\varepsilon_{n},\,f_{n}(r),\,\phi_{n}(r)\}_{n=0}^{\infty}$, with excitation index $n\in\mathbb{N}$ (see Lions [7]). For each eigenfunction $f_{n}(r)$, we associate the velocity field $v_{n}(r)$ defined in (2), that represents the analytical counterpart of experimental rotation curves (in the definition, we omit the factor $4\pi$ from angular integration for simplicity).
@@ 2.1 Implementation
Eigenstates $\{\varepsilon_{n},\,f_{n}(r),\,\phi_{n}(r),v_{n}(r)\}_{n=0}^{\infty}$ are numerically simulated up to high excitation indices ($n\leq 80$) following the implementation described by Bernstein et al. [14] (see also [16]). The numerical scheme employs two nested iterative procedures: an outer iteration handles the representation of the infinite domain, while an inner iteration solves problem (4c) on a fixed domain. Specifically, the outer iterative procedure progressively extends the computational domain until the eigenvalue correction falls below a specific tolerance. The inner procedure decouples problem (4c) by using previous-step approximations (or an initial guess for the first step) for the unknowns $f(r)$ and $\phi(r)$, solving the resulting approximate problem to achieve increasingly accurate solutions.
To verify algorithmic robustness, simulations were performed with various spatial grid refinements. This validation process allowed to monitor the quality of the results, confirming reliability up to $n=80$, while results for higher excitation indices were deemed insufficiently accurate for this study and excluded from the analysis.
@@ 3.1 Eigenfunctions: main features and notation
When applying these eigenstates to model galactic dark matter clusters, the squared modulus $|f_{n}|^{2}(r)$ represents the dark matter density distribution. Its three-dimensional structure, revealed through the simulations and shown in Figure 1, describes dark matter as arranged in concentric spherical shells of decreasing density. The quantitative study that follows characterizes this structure by deriving precise heuristic laws, that describe its dependence on the excitation index $n$.
@@ 3.5 Eigenvelocities
In the large $n$ limit the slopes approach zero, consistent with the experimental observations of flat plateaux.
@@
where $\tilde{v}_{2n}(n)$ is obtained by combining Equations (12a) and (12b), $\tilde{v}_{2n}(n)\equiv\tilde{v}_{2n}(\tilde{r}_{2n}(n))$. Figure 13 reports eigenvelocities rescaled according to Equation (13). The plot shows how the simulated rotation curves, originally shown in Figure10, collapse onto a single average curve after rescaling, revealing a universal behavior. This finding aligns well with physical expectations (see e.g. Rubin et al. [25], Burstein and Rubin [26]), which typically predict common underlying features in rotation curves across different galaxies.
@@ 4 Discussion and Conclusions
This work provides a characterization of excited Schrödinger-Poisson eigenstates, extending current knowledge of their fundamental properties, with particular emphasis on their agreement with physical expectations for galactic rotation curves. The analysis is based on numerically derived heuristic laws, that describe how key properties of the eigenstates scale with the excitation index $n$.
@@
The eigenfunctions are characterized by two main components: oscillatory behavior and amplitude modulation. Unlike standard oscillatory functions, the nodal spacing increases with radial position, showing more pronounced growth at larger radii. The amplitudes decay following a power law, with deviations observed only at the last local extrema, and an exponent approaching $-1$ in the large $n$ limit. This description both aligns with and refines the widely used approximation $f_{n}(r)\sim\sin(r)r^{-1}$. The heuristic laws also provide insight into the extent of dark matter distributions, with a support – approximately defined by the radial position of the last local extremum – showing a parabolic dependence on $n$.
@@
The resulting eigenvelocity characterization aligns well with physical expectations: the mid-range oscillatory region exhibits a linear trend with a positive slope, that approaches zero in the large $n$ limit. The clear dependence on $n$ displayed by the radial position and velocity of the last extremum provide natural scaling laws, through which a universal behavior emerges for the eigenvelocities. This universality agrees with the physical expectation of common underlying properties in dark matter distributions across different galaxies.
A natural extension of this study would be the direct comparison between model predictions and observational data, which, however, requires incorporating the baryonic component into the model (see Ji and Sin [6]). A systematic investigation of this aspect is reserved for future work.

### v2

@@ Abstract
This work provides a comprehensive numerical characterization of the excited spherically symmetric stationary states of the Schrödinger-Poisson problem. Through numerical computation of highly excited eigenstates, novel heuristic laws are proposed, which describe how their fundamental features scale with the excitation index $n$. Key characteristics of the eigenfunctions include: the effective support, which exhibits a parabolic dependence on the excitation index; the distances between adjacent nodes, whose pattern varies regularly with $n$; and the oscillation amplitude, which follows a power law with an exponent approaching $-1$ for large $n$. Based on the eigenfunctions, eigenvelocities are conveniently defined. They exhibit a mid-range oscillatory region with an average linear trend, whose slope approaches zero in the large $n$ limit; and they are characterized by heuristic scaling relationships with the excitation index $n$, revealing an intrinsic universal behavior.
@@
Schrödinger-Poisson , Scaling , Excited stationary states , Rotation curves , Universality
@@
[1] organization=Department of Mathematics “Tullio Levi-Civita”, University of Padova, addressline=Via Trieste, 63, city=Padova, postcode=35131, country=Italy
@@
[3] organization=Padua Quantum Technologies Research Center (QTech), addressline=Via Gradenigo, 6/A, city=Padova, postcode=35131, country=Italy
@@ 1 Introduction
The Schrödinger-Poisson model, or Schrödinger-Newton model, embeds a nonlinear, nonlocal effect in a Schrödinger dynamics through the coupling with a Newtonian potential. It describes the evolution of a scalar matter field $\psi(t,x)$ via the Schrödinger equation, where $|\psi|^{2}(t,x)$ represents the matter density distribution of the field. The matter field is subjected to self-gravity, which is modeled through a Poisson potential $\phi(t,x)$, with the density $|\psi|^{2}(t,x)$ acting as the source. In dimensionless form, the system reads:
@@
This system is relevant in a wide variety of fields [1]. It was first introduced by Ruffini and Bonazzola [2] as the non-relativistic approximation of general relativistic, self-gravitating systems of bosons. It was subsequently applied in astrophysics and cosmology to study non-relativistic boson stars [3] or to model ultralight scalar field dark matter ([4, 5]). Interpreting the quantum nature of the system at a more fundamental level, Diósi [6] and Penrose and Marcer [7] used the Schrödinger-Poisson model to explore the effects of self-gravitation on a single particle wavefunction, while a more recent research line [8, 9] investigates the problem of quantum gravity, showing that a Schrödinger-Poisson nonlinearity emerges naturally from a fundamentally semi-classical interpretation of gravity. Interestingly enough, the implications of this approach might also be tested through optomechanics [10, 11] or through dilute cold atom BECs experiments [12], where isotropic condensed configurations can be obtained through an electromagnetically induced gravity-like interaction [13, 14, 15]. Other fields of application for the Schrödinger-Poisson model include non-linear optics, where it can describe the propagation of light in thermo-optical media [16, 11], plasmas and electrons in semiconductors or metallic structures (see Paredes et al. [1] and references therein).
Despite the wide popularity of the Schrödinger-Poisson system, the understanding of its mathematical structure remains incomplete. In this work we focus on radially symmetric stationary states, which are solutions to system (1) of the form $(e^{i\varepsilon t}f(r),\phi(r))$. From Lieb [17] and Lions [18], they are known to comprise an infinite, discrete family $\{\varepsilon_{n},\,f_{n}(r),\,\phi_{n}(r)\}_{n=0}^{\infty}$, with positive eigenvalues $\varepsilon_{n}>0$, real eigenfunctions $f_{n}(r)$ and excitation index $n\in\mathbb{N}$. Currently, no analytic expressions exist for these eigenstates, which are therefore mainly known through numerical computations [19, 20] and asymptotic estimates [21]. As shown in Figure 1, the $n$-th excited stationary state represents matter distributions as $n+1$ concentric shells, separated by voids and exhibiting decreasing density outwards. A rigorous characterization of this structure and its dependence on the excitation index $n$ remains incomplete, which this paper aims to address.
Figure 1: Three-dimensional section of the matter distribution, as predicted by the $8$-th stationary state $f_{8}(r)$. For visualization purposes, the modulus $|f_{8}|(r)$ is shown rather than the actual predicted matter density $|f_{8}|^{2}(r)$. The structure exhibits $9$ shells of decreasing density separated by $8$ voids (shown in dark blue).
Given a predicted matter density distribution $f_{n}^{2}(r)$ , one can calculate the tangential velocity of a test object in circular orbit at radius $r$ from the distribution’s center. This calculation assumes the velocity is determined by the gravitational force exerted by matter within a sphere of radius $r$:
@@
This type of prediction, where a velocity profile is derived from the density distribution of a proposed model, is typical of galactic dark matter research, as in this context it can be compared with experimental rotation curves of spiral galaxies [19, 22, 23, 24, 25, 26], providing both model validation and deeper physical insights into the problem. The use of excited Schrödinger-Poisson stationary states as models for ultralight galactic dark matter distributions and the approximate behavior of the corresponding rotation curves were explored in a series of works [19, 27, 28], matching some relevant physical expectations [29, 30, 31]. However, their role was reconsidered after the discovery of numerical instability [32, 33] and the development of more detailed three-dimensional simulations of system (1) [34]. The latter showed the formation of a flat, oscillating core surrounded by an envelope of highly energetic granular structures, which continuously fluctuate in density and velocity. The solid angle averages and the time averages of the resulting anisotropic, time-dependent density configurations provide a universal density profile, whose core is well fitted by the ground state of the Schrödinger-Poisson problem (1) and whose halo resembles the Navarro-Frenk-White profile of standard CDM [34, 35]. The subsequent literature aimed to construct on-demand density profiles describing such average behaviors and to study their stability with simulations [35], eventually fitting them to experimental data [22, 23]. Such effective density profiles usually involve the ground state of the Schrödinger-Poisson problem (1), which behaves as a stable attractor [33] (see also [35] for a three-dimensional generalization), while a wider variety of choices are adopted to model the average behavior of the granular envelope. The excited Schrödinger-Poisson stationary states are still used in some contemporary descriptions of dark matter density profiles, which assume simultaneous occupation of several states resulting in stable effective multimodal configurations [36, 37, 35]. Despite the nonlinearity of the problem, which complicates the use of knowledge on excited stationary states to study superposed configurations, we believe that a refined understanding of the properties of such states may be of interest, independently of the specificities of the possible applications. Let us also mention that, even though the excited Schrödinger-Poisson eigenstates are unstable, some recent results showed that the highly excited ones have a lifetime that increases with $n$ [21] which, in view of the broad range of physical applications in which the Schrödinger-Poisson model is applied, opens the possibility for further analysis on this topic. A comprehensive study on the stability issue, whose interest goes beyond the specific application to the dark matter problem, is out of the scope of this work and we just include some remarks in the conclusive Section of this paper, postponing a more thorough discussion to future works.
Building on these considerations, we include in our study a comprehensive, quantitative description of the eigenvelocities (2) arising from the Schrödinger-Poisson model, characterizing their key features as explicit functions of the excitation index $n$.
@@ 1.1 Comparison with previous works and main results
However, a comprehensive analysis of how properties scale with the excitation index $n$ remains lacking. To address this gap, we propose quantitative, heuristic laws derived from numerically computed highly excited stationary states ($n\leq 80$), aiming to provide foundations for future analytical and physical studies. Specifically, we analyze the eigenfunctions’ structure, characterizing them as oscillatory functions modulated by a power law. The radial position of the last oscillation, which approximately defines the eigenfunction support, exhibits a clear parabolic dependence on $n$, characteristic of Keplerian problems like the Bohr model for the hydrogen atom (see e.g. [45]). The spacing between adjacent nodes increases with the radial position, following a consistent pattern across increasing $n$ values – showing modest increases at small radii and more pronounced increases at larger radii. This behavior constitutes a key correction to the $\sin{(r)}r^{-1}$ approximation, which fails to account for irregular nodal spacing. Finally, the power law governing amplitude modulation is characterized by fitting local extrema at small radial positions, with deviations observed near the last extremum. The power law exponent approaches $-1$ asymptotically as $n$ increases, recovering the $r^{-1}$ approximation in the large $n$ limit.
From a physical perspective, rotation curves are used in the context of galactic dark matter distributions as the primary experimental reference to compare with analytical predictions. In this framework, observed rotation curves typically show mid-range oscillatory behavior with an approximately linear trend, forming a plateau. This plateau generally exhibits a near-flat slope, although both positive and negative slopes have also been documented (see e.g. [30, 46, 22]). Moreover, physical intuition suggests that galactic-scale dark matter clusters, which generate these rotation curves, might exhibit some kind of universal behavior [31, 29], possibly in an average sense [35]. Sin [19] first noted that the key features displayed by the eigenvelocities in the Schrödinger-Poisson model align well with these physical observations, especially when taking into account the presence of a baryonic bulge [27]. As already discussed, the interpretation of these similarities is to be intended with the due caution. In view of the possible role of the excited stationary states in contemporary models, as the multimodal ones, we still add some comments, refining the observations on the mathematical structure of the eigenvelocities and emphasizing in particular the large-$n$ asymptotic behaviors, for which the instability issue might have a reduced impact [21]. We note that the large-$n$ approximation $f_{n}(r)\sim r^{-1}$ derived from the local extrema fit implies an approximately constant eigenvelocity (2), consistent with an expected flat plateau. Our systematic results refine this intuition, showing how the slopes of the mid-range eigenvelocities decay with $n$ according to a power law. In addition, the velocity at the last local oscillation exhibits a clear dependence on its radial position, which itself varies with $n$. This relationship reveals appropriate scaling laws under which all predicted eigenvelocities converge to a similar shape, independent of $n$, exhibiting an intrinsic universal behavior that might be useful in modeling multimodal configurations with average common features.
@@ 2 Model
The Schrödinger-Poisson model (1) describes the dynamics of a density $|\psi|^{2}(t,x)$ under self-gravity. The model couples a Schrödinger equation for the matter field $\psi(t,x)$ with a Poisson equation for the potential $\phi(t,x)$. Its stationary states are expressed in the form $(e^{i\varepsilon t}f(x),\phi(x))$, where the phase oscillation in time naturally emerges from the norm conservation law $\partial_{t}(\int_{\mathbb{R}^{3}}|\psi(t,x)|^{2}\mathop{}\!\mathrm{d}^{3}x)=0$, with $\varepsilon$ representing the associated energy parameter. The spatial component $f(x)$ of the matter field can be assumed real without loss of generality.
@@
where equation (4a) is known as the Choquard equation and represents a nonlinear eigenvalue problem, with eigenvalue $\varepsilon$ and eigenfunction $f(x)$. Following the prevailing approach in the literature, we focus on spherically symmetric solutions, which form an infinite, discrete family $\{\varepsilon_{n},\,f_{n}(r),\,\phi_{n}(r)\}_{n=0}^{\infty}$, with excitation index $n\in\mathbb{N}$ [18]. For each eigenfunction $f_{n}(r)$, we associate the velocity field $v_{n}(r)$ defined in (2), which we term eigenvelocity (in the definition, we omit the factor $4\pi$ from angular integration for simplicity).
@@ 2.1 Implementation
Eigenstates $\{\varepsilon_{n},\,f_{n}(r),\,\phi_{n}(r),v_{n}(r)\}_{n=0}^{\infty}$ are numerically computed up to high excitation indices ($n\leq 80$) following the implementation described by Bernstein et al. [20] (see also [39]). The numerical scheme employs two nested iterative procedures: an outer iteration handles the representation of the infinite domain, while an inner iteration solves problem (4c) on a fixed domain. Specifically, the outer iterative procedure progressively extends the computational domain until the eigenvalue correction falls below a specific tolerance. The inner procedure decouples problem (4c) by using previous-step approximations (or an initial guess for the first step) for the unknowns $f(r)$ and $\phi(r)$, solving the resulting approximate problem to achieve increasingly accurate solutions.
To verify algorithmic robustness, numerical computations were performed with various spatial grid refinements. This validation process allowed to monitor the quality of the results, confirming reliability up to $n=80$, while results for higher excitation indices were deemed insufficiently accurate for this study and excluded from the analysis.
@@ 3.1 Eigenfunctions: main features and notation
When applying these eigenstates to model matter fields, the squared modulus $|f_{n}|^{2}(r)$ represents the matter density distribution. Its three-dimensional structure, revealed by numerical computations and shown in Figure 1, describes matter as arranged in concentric spherical shells of decreasing density. The quantitative study that follows characterizes this structure by deriving precise heuristic laws, that describe its dependence on the excitation index $n$.
@@ 3.5 Eigenvelocities
In the large $n$ limit the slopes approach zero, consistent with the experimental observations of flat plateaux and with the result we would obtain by plugging in the definition (2) the $f_{n}(r)\sim r^{-1}$ approximation obtained for the eigenfunctions.
@@
where $\tilde{v}_{2n}(n)$ is obtained by combining Equations (12a) and (12b), $\tilde{v}_{2n}(n)\equiv\tilde{v}_{2n}(\tilde{r}_{2n}(n))$. Figure 13 reports eigenvelocities rescaled according to Equation (13). The plot shows how the numerically computed rotation curves, originally shown in Figure10, collapse onto a single average curve after rescaling, revealing an intrinsic universal behavior.
@@ 4 Discussion and Conclusions
This work provides a characterization of excited Schrödinger-Poisson eigenstates, extending the current knowledge of their fundamental structure. The analysis is based on numerically derived heuristic laws that describe how key properties of the eigenstates scale with the excitation index $n$.
@@
The eigenfunctions are characterized by two main components: oscillatory behavior and amplitude modulation. Unlike standard oscillatory functions, the nodal spacing increases with radial position, showing more pronounced growth at larger radii. The amplitudes decay following a power law, with deviations observed only at the last local extrema, and an exponent approaching $-1$ in the large $n$ limit. This description both aligns with and refines the widely used approximation $f_{n}(r)\sim\sin(r)r^{-1}$. The heuristic laws also provide insight into the extent of modeled matter distributions, with a support – approximately defined by the radial position of the last local extremum – showing a parabolic dependence on $n$.
@@
The resulting eigenvelocity characterization confirms and extends the results previously derived in the literature [19]. The mid-range oscillatory region exhibits a linear trend with a positive slope that decays with $n$ as a power law, approaching zero in the large $n$ limit. The clear dependence on $n$ displayed by the radial position and velocity of the last extremum provides natural scaling laws, through which a universal behavior emerges for the eigenvelocities, intrinsically characterizing the model.
Concerning the instability issue mentioned in the introduction, we emphasize that the excited eigenstates we study are unstable, as proved in [33], and thus they may not be the optimal candidate to directly model galactic dark matter distributions. However, we include a remark on a mathematical peculiarity, framed in a dynamical systems perspective. We observe that the instability of a stationary state does not imply unobservability. Indeed, as is well known, a dynamical system can spend a large amount of time in the neighborhood of an unstable (hyperbolic) equilibrium before moving away from it. Regarding dark matter, this possibility has often been considered, starting with the early works of Sin [19] and Lee and Koh [28], then with Guzmán and Ureña López [32], up to more recent works such as Roque et al. [21], who observed that “the shortest living unstable mode for each $n$ (i.e., the one with the largest real part) has a lifetime that increases with $n$. In this sense, higher excited states are less unstable than lower excited ones”. This comment refers to the exponential growth factor $e^{\lambda_{n}t}$ characterizing instability, with $\lambda_{n}$ the positive real part of the stability eigenvalues representing the reciprocal of the characteristic instability time $\tau=\lambda_{n}^{-1}$. From the numerical simulations reported by [21] (see Figure 4 therein), the value $\lambda_{n}$ seems to approach zero for $n\to\infty$, implying that the characteristic instability time $\tau=\lambda_{n}^{-1}$ may diverge. The larger the instability time, the more significant the probability that, starting near an excited state, a configuration close to that state can be observed. Applying this argument to support the use of single highly excited eigenstates as models for dark matter is beyond the scope of this work. This discussion simply aims to offer a more complete overview of the role of instability, so as to give due weight to the assumptions made on this regard when setting the stability requirements a model should satisfy. In a more general perspective, these considerations just imply that instability is not necessarily a limitation for physical applications. Thus, the comprehensive mathematical characterization of the excited eigenstates presented in this work can still be useful, supporting future analytical and physical developments in different areas.
Depending on the specific applications, this work could be extended in different directions, for example by including external sources in the Poisson equation [27] or by adding self-interaction effects [28]. Examining the dependence on $n$ of the key features in these new frameworks would further clarify how they connect to the mathematical structure of the problem. A systematic investigation of this aspect is reserved for future work.